### v1

# Principled Design of Interpretable Automated Scoring for Large-Scale Educational Assessments

Abstract
AI-driven automated scoring systems offer scalable and efficient means of evaluating complex student-generated responses. Yet, despite increasing demand for transparency and interpretability, the field has yet to develop a widely accepted solution for interpretable automated scoring to be used in large-scale real-world assessments. This work takes a principled approach to address this challenge. We analyze the needs and potential benefits of interpretable automated scoring for various assessment stakeholders and develop four principles of interpretability – Faithfulness, Groundedness, Traceability, and Interchangeability (FGTI) – targeted at those needs. To illustrate the feasibility of implementing these principles, we develop the AnalyticScore framework for short answer scoring as a baseline reference framework for future research. AnalyticScore operates by (1) extracting explicitly identifiable elements of the responses, (2) featurizing each response into human-interpretable values using LLMs, and (3) applying an intuitive ordinal logistic regression model for scoring. In terms of scoring accuracy, AnalyticScore outperforms many uninterpretable scoring methods, and is within only 0.06 QWK of the uninterpretable SOTA on average across 10 items from the ASAP-SAS dataset. By comparing against human annotators conducting the same featurization task, we further demonstrate that the featurization behavior of AnalyticScore aligns well with that of humans. footnotetext: * Equal Advising
1 Introduction
Accurate and credible assessment of knowledge and skills forms the basis for effective decision making in a variety of educational contexts, from student learning and instructional design to program development and policy making Berman et al. (2019). When the set of knowledge and skills to be gauged involves complex, open-ended problem-solving and communication abilities, AI-driven automated scoring systems can offer rapid, accessible, and scalable alternatives to the otherwise labor-intensive and costly process of training and deploying human scorers Foltz et al. (2020). Automated scoring systems have been increasingly adopted across various assessment contexts over the past several decades. Today’s scoring algorithms achieve acceptable levels of scoring accuracy in various areas of human learning Whitmer and Beiting-Parrish (2023, 2024).
Despite progress, automated scoring of open-ended responses has yet to reliably obtain generalizable scoring accuracy across diverse scoring contexts. Even when automated scoring meets acceptable levels of scoring accuracy, errors or biases inherent in the scoring algorithm can profoundly harm student learning and equity, policy evaluation, and public trust Pellegrino (2022), Berman et al. (2019). For these reasons, improving transparency and interpretability in automated scoring has now become a moral imperative, not a mere technical preference Khosravi et al. (2022), Holmes et al. (2022), Memarian and Doleck (2023), Schlippe et al. (2022). Yet, in spite of the growing research on interpretable and explainable AI as well as its applications specifically within educational assessment, interpretable automated scoring remains mostly confined to academic research with limited adoption in large-scale, real-world assessment Institute of Education Statistics (2023), Whitmer and Beiting-Parrish (2023, 2024).
In this paper, we take a principled approach towards building a practical interpretable automated scoring solution for large-scale assessments. An effective interpretability solution begins by identifying the diverse needs of each stakeholder in understanding the system’s decisions, and by grounding the development of interpretable AI systems in those needs Bhatt et al. (2020), Preece et al. (2018), Páez (2019). Research on explainable automated scoring, on the other hand, has largely ignored this need-finding process. As we observe later in this paper (Section 2), this neglect has often led to several claimed interpretability solutions that fail to address the diverse and nuanced interpretability needs of the human actors in educational assessment.
We identify the needs and benefits of model explanations for various large-scale assessment stakeholder groups consisting of test takers, assessment developers, and test users (Section 2.1). Targeted at those needs, we develop the principles of faithful, grounded, traceable, and interchangeable model interpretations for AI-driven automated scoring (Section 2.2).
We further illustrate the feasibility of implementing these principles in practice and establish a concrete baseline for future work (Section 3). AnalyticScore is the first interpretable automated short-answer scoring framework to embody our principles. It operates by extracting explicitly identifiable elements from unannotated response texts and featurizing each response into human-interpretable values based on those elements. These features are input to an intuitive ordinal logistic regression module for scoring.
We measure the performance of AnalyticScore on a real-world response dataset by measuring (1) scoring accuracy and (2) alignment of featurization behaviors with human judgments (Sections 4 and 5). AnalyticScore outperforms many uninterpretable scoring methods, and is within only 0.06 QWK of the uninterpretable SOTA on average across 10 items from the ASAP-SAS dataset. The featurization behavior of AnalyticScore also aligns well with humans (0.90, 0.72, 0.81 QWK across assessment areas). Our findings indicate strong potential for implementing accurate and well-aligned interpretability solutions that meet the real needs of assessment stakeholders.
Automated Scoring and Interpretable AI
As AI-driven automated scoring systems became more complex and opaque, researchers have increasingly noted the need to enhance the transparency of these systems through model explanations Bennett and Zhang (2015), Bauer and Zapata-Rivera (2020), Schlippe et al. (2022). Several approaches have been proposed to address interpretable automated scoring, and we discuss them in detail in Section 2.2 in connection with our four principles. Despite growing research interests, interpretable automated scoring still lacks practical adoption and meaningful field use. The 2023 NAEP Math Automated Scoring Challenge [^1] https://github.com/NAEP-AS-Challenge/math-prediction for open-ended math responses organized by the US National Center for Education Statistics (NCES) found that none of the submissions met the criteria for “interpretability” despite several methods achieving near-human scoring accuracy Institute of Education Statistics (2023), Whitmer and Beiting-Parrish (2024). This gap highlights the need for human-centered explainability solutions driven by assessment stakeholders’ real needs.
2 Building the Principles of Interpretable Automated Scoring
Insights derived from scoring support various stakeholders throughout the overall assessment process. Below we analyze three main stakeholder groups in large-scale assessment – test takers, assessment developers, and test users Berman et al. (2019), AERA et al. (2014). Each stakeholder group’s distinct roles and priorities uniquely shape how interpretable automated scoring can address their specific needs and improve their assessment experience.
2.1 Explainability Needs and Potential Benefits
Test Takers
The needs and benefits of interpretable scoring vary depending on the assessment type: summative or formative. Most large-scale assessments are summative assessments, which are assessments of learning that support evaluating learner achievement, assigning grades, or determining proficiency levels Harlen (2005). Because these assessments often drive high-stakes decisions, test takers need to trust the fairness and justifiability of scoring decisions Williamson et al. (2012). Provided that the scoring algorithm implements sound scoring logic, allowing test takers or their representatives to examine traceable explanations for scoring decisions can foster trust Bauer and Zapata-Rivera (2020), Ferrara and Qunbar (2022). These explanations can also support a streamlined quality control process by facilitating the identification and correction of errors, improving the overall integrity of the assessment (see Bennett and Zhang (2015) and Ferrara and Qunbar (2022)).
Formative assessments are assessments for learning, intended to guide and improve learner performance through frequent practice, progress monitoring, and skill diagnosis Wiliam (2011), Black and Wiliam (1998). In this context, the function of automated scoring is primarily to provide timely, effective and actionable feedback to support learner learning Bennett (2006), DiCerbo et al. (2020). Effective feedback should help learners understand the discrepancy between their work and a desired outcome Schwartz et al. (2016). A step-by-step explanation of the features observed in a learner’s work, coupled with human understandable descriptions of how those features were processed can be used to provide such elaborative feedback.
Assessment Developers
Scoring algorithms should reliably identify evidence of the constructs (target knowledge, skills, and abilities) measured by the task Bejar et al. (2016). Understanding the types of evidence that an automated scoring algorithm reliably detects also informs other key aspects of assessment design, such as construct selection and task design Bennett and Bejar (1998). Model explanations can facilitate this understanding by transparently revealing the features used by the scoring algorithm and its intermediate reasoning steps. Explanations can also help determine which parts of the algorithm can be reused, avoiding the costly and time-consuming process of training a new scoring algorithm for each new task (see DiCerbo et al. (2020)).
Model explanations also yield specific insights into areas where the scoring model can be improved and how. Scoring models often need to be tuned for various reasons. For instance, models trained on data may reflect biases related to response strategies specific to student groups Ferrara and Qunbar (2022), Rupp (2018). Scoring models may also become less stable over time as the test-taker population and/or scoring criteria change Bejar et al. (2016). Transparent inspection of model decisions helps identify problematic model elements, enabling targeted data collection and modified training objectives to improve the model.
Test Users
Test users, including professionals who select and administer tests, educators, administrators, and policymakers, depend on score reliability and interpretation validity to make system-level decisions or instructional differentiation. Their reliance on the integrity and validity of scores to drive decisions is significant AERA et al. (2014). Model explanations provide concrete evidence to validate the choice of the scoring model [^2] More examples of validity arguments on the use of automated scoring can be found in (Bennett and Zhang, 2015, Table 7.7).. This includes understanding whether the extracted features and scoring logic fully capture the rubric and the construct definition, and whether the internal structure of the automated scores align with the construct of interest Bennett and Zhang (2015).
2.2 The FGTI Interpretability Principles
We develop four foundational interpretability principles – Faithful, Grounded, Traceable, and Interchangeable (FGTI) – targeting the needs and benefits of large-scale assessment stakeholders from Section 2.1. Our first foundational principle is that explanations should be faithful Jacovi and Goldberg (2020). Faithfulness is an important requirement in many high-stakes applications of interpretable AI Rudin (2019). Similar expectations extend to assessments, and all of the needs and benefits outlined in Section 2.1 depend crucially on faithfulness.
Principle 1 (Faithful).
Explanations of scoring decisions should accurately reflect the computational mechanism behind the scoring model’s prediction.
A notable example of unfaithful scoring explanations are texts produced by prompting LLMs to generate an explanation (e.g., Lee et al. (2024) and Li et al. (2025)). Stepwise reasoning verbalized by LLM through prompting strategies such as chain-of-thought Wei et al. (2022) are not explanations of their internal computation Sarkar (2024) and often fail to reflect the model’s true reasoning behavior Turpin et al. (2023), Arcuschin et al. (2025). Moreover, LLMs are highly sensitive to superficial changes in prompts and input text, frequently exhibiting inconsistent judgments Wang et al. (2024). Therefore, even when LLMs achieve high scoring accuracy, prompting them for explanations cannot reliably address the stakeholder needs identified in Section 2.1.
Next, the model should use meaningful features that are explicitly linked to each student’s work and rely only on those features for the downstream computation.
Principle 2 (Grounded).
Initial features computed by the scoring model should represent human-understandable, explicitly identifiable elements of student work and item task.
Regardless of the routine used to derive those features, these feature values should possess meaning that is understandable to humans and be explicitly based both in the student work and item task. For instance, cosine similarity of sentence embeddings used as an input feature (e.g., Condor and Pardos (2024)) is less human-understandable than discrete features whose values are associated with clear, verbalizable meaning. Having features that are grounded addresses the need to scrutinize the features of the scoring engines.
How should the model process these features to ultimately produce a final score? Scoring is inherently an evidentiary reasoning process, where elements of the student response and item tasks serve as evidence to support the inference about student knowledge and skills that the score represents DiCerbo et al. (2020), Mislevy (2020). Stakeholders need to be able to inspect and interact with the internal structure of the scoring model to ensure soundness, construct-relevance, fairness, and model understanding (Section 2.1). To meet this need, the model’s evidentiary reasoning process must be decomposable into clear, sequential steps that a human could reliably execute and possibly intervene. Our next 2 principles state that the scoring model should be conducive to this decomposition and intervention:
Principle 3 (Traceable).
The scoring model should consist of subroutines that each represent a specific, well-defined evidentiary reasoning step on clearly specified inputs.
Principle 4 (Interchangeable).
A human should be able to act interchangeably on each of the reasoning subroutines.
Not all intermediate representations calculated by the model need to be understandable by human, but the reasoning subroutines should collectively account for the entire scoring logic. Moreover, humans should be able to act interchangeably with each decomposed module and replace the module outputs with human-generated results if deemed necessary.
Many proposed interpretability approaches are not grounded, traceable, or interchangeable in the sense described above. These include, for instance, calculating feature importance values Kumar and Boulanger (2021, 2020), Schlippe et al. (2022), Asazuma et al. (2023), displaying feature attribution maps Schlippe et al. (2022), Li et al. (2025), or presenting confidence metrics for scoring decisions Conijn et al. (2023). This limits the capacity to thoroughly inspect the model’s features and internal structure, which is critical to meeting the needs of the stakeholders.
Figure 1: Schematic of the AnalyticScore framework. The example question is: “Explain how pandas in China and koalas in Australia are similar, and how they both are different from pythons.”
3 A Principled Framework for Interpretable Automated Scoring
How would the FGTI principles be implemented in practice? To illustrate the feasibility of implementing these principles and to set a baseline for future research, we present AnalyticScore as a reference framework in the domain of short-answer scoring. In this setting, students write a short 1-5 sentence answer in response to an assessment item which is scored with an emphasis on content correctness and demonstrated reasoning Leacock and Chodorow (2003), Shermis (2015). The scoring model has access to a training set of student response texts paired with human-annotated scores $(r_{1},s_{1}),...,(r_{n},s_{n})$ and possibly additional unannotated responses $\{r_{n+1},...,r_{m}\}$. The goal at inference time is to predict the score $s$ for a new response $r$.
AnalyticScore (Figure 1) is a 3-phase, LLM-based framework grounded in our four principles of interpretable automated scoring. Phase 1 identifies explicitly grounded analytic components to be used. Phase 2 catalogs, or featurizes, the presence of these components in student responses. Phase 3 uses the features to compute a score. Phases 1 and 2 depend only on the response texts without any annotations. Human score labels are only used with Phase 3.
3.1 Phase 1: Extracting Analytic Components.
With the response texts from the training set (and optionally assessment content), AnalyticScore first extracts a set of analytic components, which are explicitly identifiable elements of student responses described in Principle 2. In this work, we consider a specific type of components which are representative, atomic units of explicit statements, arguments, or claims as in Figure 1.
$c_{1},...,c_{k}]=\text{Extract}(r_{1},...,r_{m}),$
Component extraction is implemented using an LLM with the prompts shown in Figure 2. Having too many analytic components could diminish the interchangeability (Principle 4) of the overall framework by exploding the number of features used in scoring Lipton (2018), so we limit to generating 15 components per request.
3.2 Phase 2: Featurizing Responses
Once the analytic components have been identified, student responses are featurized according to the presence of these components $c_{1},...,c_{m}$ in each response $r$. This step uses a labeling function $f(r;c)$ whose outputs are associated with human-understandable meaning (Principles  2 &  3). The exact label definitions used can be selected using natural language. In this work, we explore the following general purpose labeling function for $f(r;c)$:
[…]
We implement $f(r;c)$ using a Chain-of-Thought Wei et al. (2022) prompting template shown in Figure 2333Note that we CoT is used solely as a prompting technique, and the generated “thoughts” are explicitly discarded. . Inspired by the self-consistency decoding strategy for LLMs Wang et al. (2022), we apply the first-to-three aggregation rule to consider the possibly diverse interpretation of the labeling criteria when selecting the final output. Easily interpretable one-hot encodings of each $f(r;c_{m})$ are then concatenated to produce a $3m$-dimensional binary featurization of $r$:
$\text{Featurize}(r)=\text{OneHot}(f(r;c_{1}))\operatorname*{\mathbin{\|}}\cdots\operatorname*{\mathbin{\|}}\text{OneHot}(f(r;c_{k}))$
Figure 2: Prompts used in AnalyticScore
Distilling LLM Featurizer into Open Source
Using proprietary LLMs for featurization can quickly become too expensive in large-scale assessment settings, especially with many analytic components. To avoid the linearly growing cost of featurization, we supervised fine-tuned a small open-source model using a subsample of $(r,c)$ pairs, where $r$ is a response from the training set and $c$ is an analytic component from Phase 1. More specifically, we randomly sampled 10k pairs across all 10 items, calculated the featurization labels on these samples using o4-mini, and collected the full LLM model requests and outputs generated during this process which aligned with the aggregated final decision. This dataset was used to fine-tune Llama-3.1-8b-instruct with QLoRA Dettmers et al. (2023).
3.3 Phase 3: Logically Traceable Scoring
Based on the featurized responses, a traceable and interchangeable model (Principle 3 and 4) is selected and trained using the labeled response pairs $(r_{1},s_{1}),...,(r_{n},s_{n})$. Given the nature of the score categories, we employ the Immediate-Threshold variant of Ordinal Logistic Regression Rennie and Srebro (2005), Pedregosa et al. (2017) as our scoring module. Combined with the one-hot encoding featurization from Phase 2, the resulting algorithm calculates the sum of weights for each component and feature label: $\eta=\sum^{c}_{i=1}w_{i,f(r,c_{i})}$, where $w$ are the trained weights. Scores are determined by comparing $\eta$ to a set of learned thresholds $\theta_{j}$; the predicted score corresponds to the ordinal category $j$ for which $\theta_{j}\leq\eta<\theta_{j+1}$. $\eta$ can be understood as “evidence values” used for scoring.
3.4 Analysis of AnalyticScore
An example of AnalyticScore’s model explanation is shown in the right panel of Figure 1. By demonstrating human-understandable features of the response (Principle 2) and the exact decision process (Principle 3), the explanation transparently and faithfully reveals the actual scoring mechanism used (Principle 1). If, based on the explanation, the model is suspected to have made an error (e.g., C6 should be a check, not a triangle), a human inspector can modify the featurization and rerun the scoring algorithm (Principle 4), which is also how the “if instead…” explanation is generated.
The structure of AnalyticScore’s scoring model is akin to Concept Bottleneck Models Koh et al. (2020), Yang et al. (2023) in that we enforce a layer of intermediate representations with human-understandable “concepts.” Our framework ensures that the intermediate features have human-understandable values that are associated with explicitly identifiable elements (Principle 2), as opposed to characteristics that are inferred from the response.
4 Evaluating AnalyticScore
Having introduced AnalyticScore and discussed its interpretability, we now evaluate its scoring performance and how its featurization aligns with human judgments on a real-world response scoring dataset.
Token Len. Train Valid Test Assessment Area
Q1 $47.5\pm 22.2$ 1,341 331 557 Science
Q2 $59.2\pm 22.6$ 1,024 254 426
Q3 $47.9\pm 14.6$ 1,445 363 406
Reading
(Informational Text)
Q4 $40.3\pm 15.5$ 1,308 349 295
Q5 $25.1\pm 21.5$ 1,459 336 598 Science
Q6 $23.8\pm 22.6$ 1,418 379 599
Q7 $41.3\pm 25.1$ 1,432 367 599
Reading
Q8 $53.0\pm 32.6$ 1,446 353 599
Q9 $49.7\pm 36.3$ 1,453 345 599
Reading
(Informational Text)
Q10 $41.1\pm 28.5$ 1,314 326 546 Science
Table 1: ASAP-SAS dataset detail by item.
Dataset
The ASAP-SAS dataset Shermis (2015) [^4] https://www.kaggle.com/competitions/asap-sas/data is the largest publicly available dataset with short answer responses from schoolchildren for 10 different open-ended exam questions. Human raters double-scored and assigned a single number to each student response using a 3 or 4 point rubric. The assessment area for each question, as well as the sample sizes and response lengths are reported in Table 1. We use the original test set and split the public training set into training and validation sets with a 8:2 ratio.
AnalyticScore Implementation Details
For each assessment item, we used GPT-4.1 as the base LLM and extracted 15 analytic components except for Q7. This item uses a two-part scoring scheme to separately assess a character trait identified from the reading and its supporting evidence. We extracted 15 analytic components from each part, totaling 30 components. For the featurizer, we experimented with GPT-4.1-mini and Llama-3.1-8B-Instruct as our base LLM, each with temperature set to 0.7 and 1.0. We distilled the Llama featurizer for 2 epochs using a batch size of 4 and learning rate of 1e-4. All model calls were made through the official OpenAI API. Fine-tuning was conducted on an Ubuntu 20.04 machine with 2 RTX A6000 GPUs (49Gb memory), 16 AMD EPYC 9224 24-Core Processors, and 250Gb of CPU RAM.
4.1 Scoring Accuracy Experiment
We measured scoring accuracy in terms of quadratic weighted kappa (QWK) against the model scores in the test set, following the convention of the automated scoring literature Shermis (2015), Institute of Education Statistics (2023).
The following baseline models were compared :
Few-Shot Prompting:
We few-shot prompt GPT-4.1 with 10 randomly selected responses from each score category, including a rubric for the score categories.
Supervised Fine-tuned LLM:
The following LLM-based classifiers were fine-tuned on the response-score pairs: BERT Devlin et al. (2019), DeBERTa He et al. (2020), Llama-3.1-8b, and Llama-3.1-8b-Instruct Grattafiori et al. (2024). We also fine-tune Llama-3.1-8B-Instruct with a rubric of the score categories added to the input.
Automated Scorer Baselines:
Methods included are: AutoSAS Kumar et al. (2019), AsRRN Li et al. (2023), and NAM Condor and Pardos (2024).
The only baseline method that has aspects of interpretability is NAM. This method requires hand-crafting a specific form of rubric describing the key phrases and concepts to be used by the response. Using sentence embeddings with n-gram matching as its features, this method implements a logistic regression score classifier. To implement this baseline, we replace the rubrics with the analytic components extracted by our AnalyticScore.
4.2 Featurization Alignment Experiment
The feature labeling task described in Figure 2 was designed to produce human-understandable features (Principle 2). But how well does the LLM’s featurization behavior align with how humans actually understand this task? Even more fundamentally, how well do humans themselves agree in their understanding of this task?
To answer these questions, we sampled 50 (response, analytic component) pairs for each of the 3 assessment areas. To ensure balanced representation, the sample included a balanced number of pairs from each of the three score categories, as initially determined by the GPT-4.1-mini featurizer. We then asked 7 human annotators to conduct the labeling task on these samples. The human annotators consisted of five volunteers from an R1 university and two of the study’s authors. None of the annotators had prior exposure to any of the LLM’s featurization outputs. All annotators had advanced academic training (PhD-level) and teaching experience, five of whom have been instructors at the primary, secondary, and/or post-secondary level.
The annotators received an oral presentation of the purpose of the study along with links to 3 Qualtrics forms to be filled out, one in each assessment area. The form reiterated the study’s purpose, explained the task, and presented 50 items to annotate, each containing the context of the assessment item and the same featurizer prompt shown in Figure 2. The overall process took each annotator between 2.5 to 3.5 hours.
Aggregate Human label was generated by majority voting (ties resolved randomly). We calculated inter-rater reliability among human labelers (Krippendorff’s $\alpha$) and alignment between each LLM featurizer and aggregate human labels (QWK and class-wise F1). We report the 95% bootstrap CI of each metric, reweighting the sampling probability to account for the initial balanced sampling of score categories.
5 Experiment Results
5.1 Scoring Accuracy Results
Table 2 shows the results of the scoring accuracy experiments. Across items and within each assessment area, AnalyticScore outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. several automated scoring baselines on average and, given its interpretability, achieves reasonable performance compared to state-of-the-art black-box models. Except for the untuned Llama featurizer, each AnalyticScore variant outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. the few-shot prompting and automated scoring baselines. Compared to the best-performing models in each assessment area, these three AnalyticScore models are, on average, within 0.06 QWK over all items, 0.04 QWK for Science, 0.08 QWK for Reading (Informational Text), and 0.09 QWK for Reading (Literature) items.
Also noticeable is the striking improvement111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. in the performance of the Llama featurizer post-distillation, with an average increase of 0.13 QWK. The distilled Llama featurizer performs comparably to both variants of GPT-4.1 mini. Increase in average QWK is most notable for Science items (+0.19), followed by Reading (Literature) (+0.12) and Reading (Informational Text) (+0.05).
Q1 Q2 Q3 Q4 Q5 Q6 Q7 Q8 Q9 Q10 All Avg. Sci Avg. R(Inf) Avg. R(Lit) Avg.
[…]
AnalyticGrade
[…]
Fewshot
[…]
Supervised LLM
[…]
Baseline
[…]
Table 2: Test-time Quadratic Weighted Kappa (QWK) of scoring models per item, along with average per assessment area. Best, second-best, and at human-level performance scores are marked respectively. Sci.: Science (Q1,2,5,6). R(Inf): Reading(Informational Text) (Q3,4,9). R(Lit): Reading(Literature) (Q7,8). (*) are methods that are considered interpretable.
5.2 Featurization Alignment Results
[…]
Next, alignment between each featurizing model with the human ratings using majority vote is shown in Table 3. Most notably, the distilled Llama featurizer achieves substantially high agreement with the aggregate human features across all assessment areas. Other featurizers also achieve high agreement in Science and Reading (Information Text) but achieves moderate agreement in Reading (Literature).
Assessment Area Featurizer Model QWK Label Distribution6 Label-wise F1
2 1 0 2 1 0
Science Human 15.32% 3.70% 80.98%
GPT-4.1-mini (0.89, 0.89) 7.56% 12.59% 79.85% (0.83, 0.84) (0.20, 0.22) (0.96, 0.96)
o4-mini (0.94, 0.95) 14.35% 8.17% 77.47% (0.93, 0.93) (0.49, 0.51) (0.98, 0.98)
Llama-3.1-8B (Distilled) (0.90, 0.90) 11.54% 5.27% 83.19% (0.89, 0.89) (0.20, 0.22) (0.97, 0.97)
Reading
(Informational Text)
Human 11.64% 18.53% 69.83%
GPT-4.1-mini (0.72, 0.72) 9.82% 22.35% 67.83% (0.68, 0.69) (0.54, 0.55) (0.87, 0.88)
o4-mini (0.81, 0.81) 18.30% 16.60% 65.10% (0.73, 0.74) (0.68, 0.69) (0.94, 0.94)
Llama-3.1-8B (Distilled) (0.72, 0.73) 20.51% 10.16% 69.34% (0.61, 0.62) (0.24, 0.26) (0.91, 0.91)
Reading
Human 9.48% 6.57% 83.95%
GPT-4.1-mini (0.54, 0.56) 2.60% 13.86% 83.54% (0.45, 0.47) (0.67, 0.69) (0.95, 0.95)
o4-mini (0.52, 0.54) 7.22% 6.80% 85.98% (0.50, 0.52) (0.12, 0.14) (0.92, 0.92)
Llama-3.1-8B (Distilled) (0.81, 0.81) 7.22% 7.52% 85.26% (0.83, 0.84) (0.20, 0.22) (0.92, 0.92)
Table 3: Alignment between LLM featurizers and Aggregate Human featurization obtained by majority voting for different models and assessment area. QWK and F1 values presented are 95% Bootstrap CI.
Assessment Area Krippendorff’s $\alpha$
Science $(0.718,0.723)$
Reading (Informational Text) $(0.696,0.700)$
Reading (Literature) $(0.669,0.678)$
Table 4: Inter-rater reliability among human raters for the featurization alignment experiment (95% bootstrap CI).
[…]
6 Discussions and Limitations
The principles outlined in Section 2.2 address a specific aspect of “interpretability” in the broader domain of automated scoring. While these principles are foundational to supporting the design of a valid scoring process and fostering trust in the scoring system, simply adhering to the principles does not, on its own, guarantee that the needs and potential benefits of the stakeholders (Section 2.1) will be met.
Educational assessment literature is ripe with guidelines, frameworks, and best practices for ensuring that automated scoring properly serves the stakeholders’ needs and produce valid, reliable, and fair scoring results (e.g., Bennett and Bejar (1998), Bennett and Zhang (2015), Williamson et al. (2012), Bejar et al. (2016)). These studies emphasize that evidence for the validity of automated scoring should be collected throughout the assessment process from a variety of evidentiary sources, such as model features, agreement with human raters, treatment of unusual responses, generalizability of score interpretations, population invariance of scores, and impact on teaching and learning Bennett and Zhang (2015).
7 Conclusion
AI and education research community has yet to produce a practical interpretability solution for automated scoring in large-scale educational assessments despite a pressing need. To address this challenge, we analyzed the needs and potential benefits of interpretable automated scoring for various assessment stakeholders (students, assessment developers, and test users) and developed 4 foundational principles – Faithful, Grounded, Traceable, and Interchangeable (FGTI) – aimed at addressing those needs. We also demonstrated the feasibility of implementing these principles by developing AnalyticScore for short-answer scoring. This framework generates human-interpretable features for each response based on explicitly identifiable elements, and uses an intuitive ordinal logistic regression scorer. On a real-world short-answer scoring dataset, AnalyticScore outperforms many uninterpretable scoring methods, achieves a narrow performance gap relative to the uninterpretable SOTA, and demonstrates featurization behaviors that align with human judgment. Our findings show strong promise for implementing accurate and well-aligned interpretability solutions that address the real needs of assessment stakeholders. We hope our work illuminates exciting new directions in developing practical and effective interpretable automated scoring for large-scale educational assessments.